# Regularity of vector fields with piecewise regular curl and divergence

### v2

@@ 3 Differential operators on surfaces and traces of ${\bf H}(\operatorname{curl},\Omega)$
For most of this section we assume that $\Sigma$ is a closed and orientable $\mathcal{C}^{2}$-surface consisting of $n^{\prime}\geq 1$ simply connected components [^2] You can think of $\Sigma$ being a simply connected component $\Gamma_{j}$ of $\Gamma$ or an interface component $\mathcal{I}_{j}$., and ${\bf n}:\Sigma\rightarrow\mathbb{S}_{2}$ denotes a unit vector field normal to $\Sigma$.
@@
Note that $-\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$ map scalar functions to tangent fields. Thus, their adjoint operators $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ map tangent fields to scalar functions and satisfy
@@
For simplicity we chose to introduce $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ as adjoint operators to $-\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$, respectively. For a more rigorous approach we refer e.g. to [21, Section 2.5.6]. Moreover, the definition of the above surface differential operators can be extended to Lipschitz surfaces, [2].
@@
For any $\ell\in\mathbb{N}_{0}$ and any closed and orientable $\mathcal{C}^{\ell+2}$-surface $\Sigma$ consisting of $n^{\prime}\geq 1$ simply connected components, the surface differential operators $\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$ extend to bounded linear operators $\nabla_{\Sigma}:\operatorname{H}^{\ell+1/2}(\Sigma)\rightarrow{\bf H}^{\ell-1/2}_{T}(\Sigma)$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}:\operatorname{H}^{\ell+1/2}(\Sigma)\rightarrow{\bf H}^{\ell-1/2}_{T}(\Sigma)$. Moreover, the operators $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ extend to bounded linear operators $\operatorname{div}_{\Sigma}:{\bf H}^{\ell+1/2}_{T}(\Sigma)\rightarrow\operatorname{H}^{\ell-1/2}(\Sigma)$ and $\operatorname{curl}_{\Sigma}:{\bf H}^{\ell+1/2}_{T}(\Sigma)\rightarrow\operatorname{H}^{\ell-1/2}(\Sigma)$.
@@ 4.1 Jump operators
Thus, for any $\mathcal{C}^{\ell+1}$-partition $\mathcal{G}$ we may define a jump operator $\left[\ \right]\hskip-2.84544pt:\operatorname{PH}^{\ell+1}(\mathcal{G})\rightarrow\operatorname{H}^{\ell+1/2}(\mathcal{I})$ piecewise by
@@
on every interface component $\mathcal{I}_{j}=\partial\mathcal{G}_{j_{1}}\cap\partial\mathcal{G}_{j_{2}}$, where without loss of generality $j_{1}>j_{2}$, and where $u_{j_{1}}:=u|_{\mathcal{G}_{j_{1}}}$ as well as $u_{j_{2}}:=u|_{\mathcal{G}_{j_{2}}}$. Similarly, for any given $\mathcal{C}^{\ell+2}$-partition $\mathcal{G}$ we may define normal and tangential jump operators $\llbracket\ \rrbracket^{\operatorname{n}}\hskip-2.84544pt:{\bf PH}^{\ell+1}(\mathcal{G})\rightarrow\operatorname{H}^{\ell+1/2}(\mathcal{I})$ and $\llbracket\ \rrbracket^{\operatorname{t}}\hskip-2.84544pt:{\bf PH}^{\ell+1}(\mathcal{G})\rightarrow{\bf H}^{\ell+1/2}_{T}(\mathcal{I})$ piecewise by
@@
According to Proposition 3.4 we have $\operatorname{curl}_{\Gamma}\nabla_{\Gamma}u=0$, thus $\operatorname{curl}_{\Gamma}({\bf g})=0$ is a necessary condition. Suppose $\operatorname{curl}_{\Gamma}({\bf g})=0$. The idea is to use a similar trick as in [17, Proof of Theorem 4.3]: According to Proposition 3.3 there holds $\operatorname{div}_{\Gamma}({\bf g})\in{\bf H}^{\ell-1/2}(\Gamma)$, thus, the zero mean-value solution $z$ of $\operatorname{div}_{\Gamma}\nabla_{\Gamma}z=\operatorname{div}_{\Gamma}({\bf g})$ is an element of $\operatorname{H}^{\ell+3/2}(\Gamma)$. We define $u$ as the solution of the Poisson problem
$-\operatorname{div}\boldsymbol{\varepsilon}\nabla u=$ $f\quad{\rm in}\ \Omega$
@@
Then we have $u\in\operatorname{H}^{\ell+2}(\Omega)$ and by construction $u$ satisfies $\nabla_{\Gamma}u={\bf g}$ on $\Gamma$.
Finally, note that the difference $v:=u_{1}-u_{2}$ of two solutions $u_{1}$ and $u_{2}$ satisfies $\operatorname{div}\boldsymbol{\nu}\nabla v=0$ in $\Omega$ and $\nabla_{\Gamma}v=0$ on $\Gamma$. The latter implies that $v|_{\Gamma}$ is constant, thus, due to $\operatorname{div}\boldsymbol{\varepsilon}\nabla v=0$, the function $v$ is constant in $\Omega$. That is, the solution space forms a one-dimensional affine subspace of $\operatorname{H}^{\ell+2}(\Omega)$. ∎
@@ 4.2 Piecewise Sobolev regularity of Poisson transmission problems
[…]
@@
Then, there holds $u\in\operatorname{PH}^{\ell+2}(\mathcal{G})$ and
@@
Let $\operatorname{div}\widetilde{{\bf f}}:=\operatorname{div}{\bf f}+\operatorname{div}\boldsymbol{\nu}\nabla w\in\operatorname{PH}^{\ell}(\mathcal{G})$ and $\widetilde{h}:=\llbracket\boldsymbol{\nu}\nabla u-\boldsymbol{\nu}\nabla w\rrbracket^{\operatorname{n}}\in\operatorname{H}^{\ell+1/2}(\mathcal{I})$. Then, $v:=u-w$ satisfies $-\operatorname{div}\boldsymbol{\nu}\nabla v=\operatorname{div}\widetilde{{\bf f}}$ on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, as well as $\llbracket\boldsymbol{\nu}\nabla v\rrbracket^{\operatorname{n}}=\widetilde{h}$ and $\left[v\right]=0$, and $\boldsymbol{\nu}\nabla v\cdot{\bf n}=f-\boldsymbol{\nu}\nabla w\cdot{\bf n}$ on the boundary $\Gamma$. Moreover, we note that $\left(v,1\right)_{\operatorname{L}^{2}(\Omega)}=0$.
The equation $\left[v\right]=0$ implies $v\in\operatorname{H}^{1}(\Omega)$, and partial integration on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ yields that $v$ satisfies
@@
From (4.4) and a Poincaré inequality (which is available due to $\left(v,1\right)_{\operatorname{L}^{2}(\Omega)}=0$) we infer that
$\left\|\nabla v\right\|_{{\bf L}^{2}(\Omega)}\leq C\left(\left\|\operatorname{div}{\widetilde{{\bf f}}}\right\|_{\operatorname{L}^{2}(\Omega)}+\left\|\widetilde{h}\right\|_{\operatorname{L}^{2}(\mathcal{I})}+\left\|\widetilde{f}\right\|_{\operatorname{L}^{2}(\Gamma)}\right),$ (4.6)
@@
Together with $v:=u-w$, the definitions of $\operatorname{div}\widetilde{{\bf f}}$, $\widetilde{f}$ and $\widetilde{h}$, as well as (4.3), the inequality (4.7) implies (4.2).
@@
Analogously as in step 1, one proves (4.5) for $v:=u-w$. Then, (4.8) follows from (4.5), the inequality $\left\|v\right\|_{\operatorname{H}^{1}(\Omega)}\leq\left\|u\right\|_{\operatorname{PH}^{1}(\mathcal{G})}+\left\|w\right\|_{\operatorname{PH}^{1}(\Omega)}$ and $\left\|w\right\|_{\operatorname{PH}^{\ell+2}(\mathcal{G})}\leq C\left(\left\|{\bf g}\right\|_{{\bf H}^{\ell+1/2}_{T}(\mathcal{I})}+\left\|u\right\|_{\operatorname{PH}^{1}(\Omega)}\right)$.
Step 3: The third and final step of the proof is to conclude (4.1). From (4.8) we see that it suffices to bound $\left\|\nabla u\right\|_{{\bf L}^{2}(\Omega)}$ in terms of $\operatorname{div}{\bf f}$, $f$, ${\bf g}$ and $h$ in suitable norms. By adding a suitable piecewise constant function to $u$ we can construct a function $\widetilde{u}\in\operatorname{PH}^{\ell+2}(\mathcal{G})$ with $\nabla\widetilde{u}=\nabla u$ on every subdomain $\mathcal{G}_{i}$ and $\left(\widetilde{u},1\right)_{\operatorname{L}^{2}(\Omega)}=0$, as well as $\left(\widetilde{u},1\right)_{\operatorname{L}^{2}(\mathcal{I})}=0$. According to step 1, the function $\widetilde{u}$ satisfies (4.2), which implies
@@
[…]
@@ 4.3 Helmholtz decompositions with piecewise regularity properties
Proof. According to [17, Lemma 2.7], there exists a decomposition ${\bf u}={\bf v}+\nabla\phi$, where ${\bf v}$ and $\phi$ satisfy $\left\|{\bf v}\right\|_{{\bf H}^{1}(\Omega)}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and $\left\|\phi\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf H}(\operatorname{curl},\Omega)}$. Due to $\boldsymbol{\nu}{\bf u}$ being orthogonal to all gradient fields, there holds $-\operatorname{div}\boldsymbol{\nu}\nabla\phi=\operatorname{div}\boldsymbol{\nu}{\bf v}$ on every subdomain $\mathcal{G}_{i}$, as well as $\boldsymbol{\nu}\nabla\phi\cdot{\bf n}=-\boldsymbol{\nu}{\bf v}\cdot{\bf n}$ on the boundary $\Gamma$. Moreover, due to ${\bf v}\in{\bf H}^{1}(\Omega)$ and $\boldsymbol{\nu}{\bf u}\in{\bf H}(\operatorname{div},\Omega)$ we have $\llbracket\nabla\phi\rrbracket^{\operatorname{n}}=-\llbracket\boldsymbol{\nu}{\bf v}\rrbracket^{\operatorname{n}}$ as well as $\llbracket\nabla\phi\rrbracket^{\operatorname{t}}=0$. Without loss of generality we may assume $\left(\phi,1\right)_{\operatorname{L}^{2}(\Omega)}=0$, and according to Lemma 4.2 the function $\phi$ satisfies
$\left\|\phi\right\|_{\operatorname{PH}^{2}(\mathcal{G})}\leq C\left\|{\bf v}\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
where we used that $\phi\in\operatorname{H}^{1}(\Omega)$ implies $\left(\left[\phi\right],1\right)_{\operatorname{L}^{2}(\mathcal{I})}=0$. This proves $(i)$.
In order to prove $(ii)$, let ${\bf u}\in{\bf H}(\operatorname{curl},\Omega)$ be given and let $\phi\in\operatorname{H}^{1}(\Omega)$ satisfy $\left(\phi,1\right)_{\operatorname{L}^{2}{\Omega}}=0$ and
@@
for all $\xi\in\operatorname{H}^{1}(\Omega)$. Note that the Lax-Milgram lemma asserts the existence of $\phi$. By construction, we have $\left\|\nabla\varphi\right\|_{{\bf L}^{2}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and a Poincaré inequality shows $\left\|\varphi\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$. Furthermore, we observe that ${\bf v}:={\bf u}-\nabla\varphi\in{\bf H}(\operatorname{curl},\Omega)$ satisfies $\left(\boldsymbol{\nu}{\bf v},\nabla\xi\right)_{{\bf L}^{2}(\Omega)}=0$ for all $\xi\in{\bf H}^{1}(\Omega)$, and according to statement $(i)$ there holds $\left\|{\bf v}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf v}\right\|_{{\bf L}^{2}(\Omega)}=C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$. This proves $(ii)$.
If $\boldsymbol{\nu}$ is continuously differentiable on $\overline{\Omega}$, then the proof follows the lines, except for the fact that the equation $-\operatorname{div}\boldsymbol{\nu}\nabla\phi=\operatorname{div}\boldsymbol{\nu}{\bf v}$ in $\Omega$ with $\boldsymbol{\nu}\nabla\phi\cdot{\bf n}=-\boldsymbol{\nu}{\bf v}\cdot{\bf n}$ on $\Gamma$ is a standard Neumann problem for $\phi$. ∎
@@
$\left(\boldsymbol{\nu}^{-1}\operatorname{curl}{\bf z},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=\left({\bf u},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}$
for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$. Note that Lemma 4.5 and the Lax-Milgram lemma assert the existence of ${\bf z}$. Moreover, we have $\left\|\operatorname{curl}{\bf z}\right\|_{{\bf L}^{2}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and exploiting Lemma 4.5 leads to

### v3

@@ 3 Differential operators on surfaces and traces of ${\bf H}(\operatorname{curl},\Omega)$
For most of this section we assume that $\Sigma$ is a closed (i.e., compact and without boundary) and orientable $\mathcal{C}^{2}$-surface consisting of $n^{\prime}\geq 1$ simply connected components [^2] You can think of $\Sigma$ being a simply connected component $\Gamma_{j}$ of $\Gamma$ or an interface component $\mathcal{I}_{j}$., and ${\bf n}:\Sigma\rightarrow\mathbb{S}_{2}$ denotes a unit vector field normal to $\Sigma$. Let us mention that for many statements of this section, the assumption of $\Sigma$ being $\mathcal{C}^{2}$ is stronger than necessary. Indeed, many results can be extended to the case of $\Sigma$ being only Lipschitz. However, for Lipschitz surfaces some statements become more technical as one has to be more careful in the analysis of the subsequently defined surface differential operators. Hence, in order to make things easier we consider only surfaces that are at least $\mathcal{C}^{2}$.
@@
Note that $\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$ map scalar functions to tangent fields. Thus, their adjoint operators $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ map tangent fields to scalar functions and satisfy
@@
For simplicity we chose to introduce $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ as adjoint operators to $\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$, respectively. For a more rigorous approach we refer e.g. to [21, Section 2.5.6]. Moreover, the definition of the above surface differential operators can be extended to Lipschitz surfaces, [2].
@@
For any $\ell\in\mathbb{N}$ and $s\in[-\ell,\ell+1]$ and any closed and orientable $\mathcal{C}^{\ell+1}$-surface $\Sigma$ consisting of $n^{\prime}\geq 1$ simply connected components, the surface differential operators $\nabla_{\Sigma}$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}$ extend to bounded linear operators $\nabla_{\Sigma}:\operatorname{H}^{{s}}(\Sigma)\rightarrow{\bf H}^{{s-1}}_{T}(\Sigma)$ and $\overrightarrow{\operatorname{curl}_{\Sigma}}:\operatorname{H}^{{s}}(\Sigma)\rightarrow{\bf H}^{{s-1}}_{T}(\Sigma)$. Moreover, the operators $\operatorname{div}_{\Sigma}$ and $\operatorname{curl}_{\Sigma}$ extend to bounded linear operators $\operatorname{div}_{\Sigma}:{\bf H}^{{s}}_{T}(\Sigma)\rightarrow\operatorname{H}^{{s-1}}(\Sigma)$ and $\operatorname{curl}_{\Sigma}:{\bf H}^{{s}}_{T}(\Sigma)\rightarrow\operatorname{H}^{{s-1}}(\Sigma)$.
Remark 3.4.
In the case $\ell=0$ and if $\Sigma$ is merely Lipschitz, the above statement is more involved, as the ranges of the surface differential operators cannot be characterized as easily by the ”typical” Sobolev spaces on $\Sigma$. For surfaces that are at least $\mathcal{C}^{2}$ things become much easier, since the associated normal vector fields are then at least $\mathcal{C}^{1}$ which leads to the surface differential operators behaving very nicely. Nevertheless we stress that the regularity assumptions on $\Sigma$ from Proposition 3.3 are not optimal and can probably be relaxed by a more thorough analysis.
Similarly, the assumption that $\Sigma$ is $\mathcal{C}^{2}$ from Lemma 3.2 and Proposition 3.5 below is probably stronger than necessary. However, for our purposes the assumption that $\Sigma$ is $\mathcal{C}^{2}$ is sufficient, hence we do not aim to extend these statements to less regular surfaces.
@@ 4.1 Jump operators
Thus, for any $\mathcal{C}^{\ell+1}$-partition $\mathcal{G}$ we may define a jump operator $\left[\ \right]\thinspace:\operatorname{PH}^{\ell+1}(\mathcal{G})\rightarrow\operatorname{H}^{\ell+1/2}(\mathcal{I})$ piecewise by
@@
on every interface component $\mathcal{I}_{j}=\partial\mathcal{G}_{j_{1}}\cap\partial\mathcal{G}_{j_{2}}$, where without loss of generality $j_{1}>j_{2}$, and where $u_{j_{1}}:=u|_{\mathcal{G}_{j_{1}}}$ as well as $u_{j_{2}}:=u|_{\mathcal{G}_{j_{2}}}$. Similarly, for any given $\mathcal{C}^{\ell+2}$-partition $\mathcal{G}$ we may define normal and tangential jump operators $\llbracket\ \rrbracket^{\operatorname{n}}\thinspace:{\bf PH}^{\ell+1}(\mathcal{G})\rightarrow\operatorname{H}^{\ell+1/2}(\mathcal{I})$ and $\llbracket\ \rrbracket^{\operatorname{t}}\thinspace:{\bf PH}^{\ell+1}(\mathcal{G})\rightarrow{\bf H}^{\ell+1/2}_{T}(\mathcal{I})$ piecewise by
@@
According to Proposition 3.5 we have $\operatorname{curl}_{\Gamma}\nabla_{\Gamma}u=0$, thus $\operatorname{curl}_{\Gamma}({\bf g})=0$ is a necessary condition. Suppose $\operatorname{curl}_{\Gamma}({\bf g})=0$. The idea is to use a similar trick as in [17, Proof of Theorem 4.3]: According to Proposition 3.3 there holds $\operatorname{div}_{\Gamma}({\bf g})\in\operatorname{H}^{\ell-1/2}(\Gamma)$, thus the zero mean-value solution $z$ of $\operatorname{div}_{\Gamma}\nabla_{\Gamma}z=\operatorname{div}_{\Gamma}({\bf g})$ is an element of $\operatorname{H}^{\ell+3/2}(\Gamma)$. We define $u$ as the solution of the Poisson problem
$-\operatorname{div}{\boldsymbol{\nu}}\nabla u=$ $f\quad{\rm in}\ \Omega$
@@
Then we have $u\in\operatorname{H}^{\ell+2}(\Omega)$ and from $\operatorname{curl}_{\Gamma}\left({\bf g}\right)=0$ and Proposition 3.5 we infer ${\bf g}=\nabla_{\Gamma}\xi$ for some $\xi\in\operatorname{H}^{1}(\Gamma)$ with zero mean. By construction of $u$ this leads to $\operatorname{div}_{\Gamma}\nabla_{\Gamma}u=\operatorname{div}_{\Gamma}\nabla_{\Gamma}z=\operatorname{div}_{\Gamma}\nabla_{\Gamma}\xi$ and by uniqueness of the zero-mean solution of this elliptic problem we infer $\nabla_{\Gamma}u=\nabla_{\Gamma}\xi={\bf g}$ on $\Gamma$.
Finally, note that the difference $v:=u_{1}-u_{2}$ of two solutions $u_{1}$ and $u_{2}$ satisfies $\operatorname{div}\boldsymbol{\nu}\nabla v=0$ in $\Omega$ and $\nabla_{\Gamma}v=0$ on $\Gamma$. The latter implies that $v|_{\Gamma}$ is constant, thus, due to $\operatorname{div}{\boldsymbol{\nu}}\nabla v=0$, the function $v$ is constant in $\Omega$. That is, the solution space forms a one-dimensional affine subspace of $\operatorname{H}^{\ell+2}(\Omega)$. ∎
@@ 4.2 Piecewise Sobolev regularity of Poisson transmission problems
$\begin{split}-\operatorname{div}\boldsymbol{\nu}\nabla u=\operatorname{div}{\bf f}&\qquad\rm{in}\ \mathcal{G}_{1}\cup\ldots\cup\mathcal{G}_{n},\\ \boldsymbol{\nu}\nabla u\cdot{\bf n}=f&\qquad\rm{on}\ \Gamma,\\ \llbracket\nabla u\rrbracket^{\operatorname{t}}={\bf g}&\qquad\rm{on}\ \mathcal{I},\\ \llbracket\boldsymbol{\nu}\nabla u\rrbracket^{\operatorname{n}}=h&\qquad\rm{on}\ \mathcal{I},\end{split}$
@@
Then, there holds $u\in\operatorname{PH}^{\ell+2}(\mathcal{G})$ and 333 By a slight abuse of notation we write $\operatorname{div}{\bf f}$ for the piecewise divergence of $\bf f$}.
@@
Let $\operatorname{div}\widetilde{{\bf f}}:=\operatorname{div}{\bf f}+\operatorname{div}\boldsymbol{\nu}\nabla w\in\operatorname{PH}^{\ell}(\mathcal{G})$ and $\widetilde{h}:=\llbracket\boldsymbol{\nu}\nabla u-\boldsymbol{\nu}\nabla w\rrbracket^{\operatorname{n}}\in\operatorname{H}^{\ell+1/2}(\mathcal{I})$. Then, v:=uwv:=u-w satisfies $\operatorname{div}\boldsymbol{\nu}\nabla v=\operatorname{div}\widetilde{{\bf f}}$ on all subdomains $\mathcal{G}_{j}$, as well as $\llbracket\boldsymbol{\nu}\nabla v\rrbracket^{\operatorname{n}}=\widetilde{h}$ and $\left[v\right]=0$, and $\boldsymbol{\nu}\nabla v\cdot{\bf n}=f-\boldsymbol{\nu}\nabla w\cdot{\bf n}$ on the boundary $\Gamma$. Moreover, we note that $\left(v,1\right)_{\operatorname{L}^{2}(\Omega)}=0$.
The equation $\left[v\right]=0$ implies $v\in\operatorname{H}^{1}(\Omega)$, and partial integration on all subdomains $\mathcal{G}_{j}$ yields that $v$ satisfies
@@
By choosing $\xi=v$ in (4.4), a Poincaré inequality (which is available due to $\left(v,1\right)_{\operatorname{L}^{2}(\Omega)}=0$) and some straightforward estimates we obtain
$\left\|\nabla v\right\|_{{\bf L}^{2}(\Omega)}\leq C\left\|\widetilde{h}\right\|_{\operatorname{L}^{2}(\mathcal{I})}+\left\|\widetilde{f}\right\|_{\operatorname{L}^{2}(\Gamma)}+\left\|\operatorname{div}\widetilde{{\bf f}}\right\|_{\operatorname{L}^{2}(\Omega)}$
for a constant $C>0$ depending only on $\boldsymbol{\nu}$ and $\Omega$, where we slightly abused notation and wrote $\operatorname{div}\widetilde{{\bf f}}$ for the piecewise divergence of $\widetilde{{\bf f}}$.Exploiting the Poincaré inequality a second time yields
$\left\|v\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|\widetilde{h}\right\|_{\operatorname{L}^{2}(\mathcal{I})}+\left\|\widetilde{f}\right\|_{\operatorname{L}^{2}(\Gamma)}+\left\|\operatorname{div}\widetilde{{\bf f}}\right\|_{\operatorname{L}^{2}(\Omega)},$
@@
Together with v:=uwv:=u-w, the definitions of $\operatorname{div}\widetilde{{\bf f}}$, $\widetilde{f}$ and $\widetilde{h}$, as well as (4.3), the inequality (4.6) implies (4.2).
@@
Analogously as in step 1, one proves (4.5) for v:=uwv:=u-w. Then, (4.7) follows from (4.5), the inequality $\left\|v\right\|_{\operatorname{H}^{1}(\Omega)}\leq\left\|u\right\|_{\operatorname{PH}^{1}(\mathcal{G})}+\left\|w\right\|_{\operatorname{PH}^{1}(\Omega)}$ and $\left\|w\right\|_{\operatorname{PH}^{\ell+2}(\mathcal{G})}\leq C\left(\left\|{\bf g}\right\|_{{\bf H}^{\ell+1/2}_{T}(\mathcal{I})}+\left\|u\right\|_{\operatorname{PH}^{1}(\Omega)}\right)$.
Step 3: The third and final step of the proof is to conclude (4.1). From (4.7) we see that it suffices to bound $\left\|\nabla u\right\|_{{\bf L}^{2}(\Omega)}$ in terms of $\operatorname{div}{\bf f}$, $f$, $\bf g$} and $h$ in suitable norms. By adding a suitable piecewise constant function to $u$ we can construct a function $\widetilde{u}\in\operatorname{PH}^{\ell+2}(\mathcal{G})$ with $\nabla\widetilde{u}=\nabla u$ on every subdomain $\mathcal{G}_{i}$ and $\left(\widetilde{u},1\right)_{\operatorname{L}^{2}(\Omega)}=0$, as well as $\left({\left[\widetilde{u}\right]},1\right)_{\operatorname{L}^{2}(\mathcal{I})}=0$. According to step 1, the function $\widetilde{u}$ satisfies (4.2), which implies
@@
$\begin{split}-\operatorname{div}\boldsymbol{\nu}\nabla u=\operatorname{div}{\bf f}&\qquad\rm{in}\ \mathcal{G}_{1}\cup\ldots\cup\mathcal{G}_{n},\\ u=f&\qquad\rm{on}\ \Gamma,\\ \llbracket\nabla u\rrbracket^{\operatorname{t}}={\bf g}&\qquad\rm{on}\ \mathcal{I},\\ \llbracket\boldsymbol{\nu}\nabla u\rrbracket^{\operatorname{n}}=h&\qquad\rm{on}\ \mathcal{I},\end{split}$
@@ 4.3 Helmholtz decompositions with piecewise regularity properties
Proof. According to [17, Lemma 2.7], there exists a decomposition ${\bf u}={\bf v}+\nabla{\varphi}$, where $\bf v$} and $\varphi$} satisfy $\left\|{\bf v}\right\|_{{\bf H}^{1}(\Omega)}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and $\left\|{\varphi}\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf H}(\operatorname{curl},\Omega)}$. Due to $\boldsymbol{\nu}{\bf u}$ being orthogonal to all gradient fields, there holds $\operatorname{div}\boldsymbol{\nu}\nabla{\varphi}=\operatorname{div}\boldsymbol{\nu}{\bf v}$ on every subdomain $\mathcal{G}_{i}$, as well as $\boldsymbol{\nu}\nabla{\varphi}\cdot{\bf n}=-\boldsymbol{\nu}{\bf v}\cdot{\bf n}$ on the boundary $\Gamma$. Moreover, due to ${\bf v}\in{\bf H}^{1}(\Omega)$ and $\boldsymbol{\nu}{\bf u}\in{\bf H}(\operatorname{div},\Omega)$ we have $\llbracket\nabla{\varphi}\rrbracket^{\operatorname{n}}=-\llbracket\boldsymbol{\nu}{\bf v}\rrbracket^{\operatorname{n}}$ as well as $\llbracket\nabla{\varphi}\rrbracket^{\operatorname{t}}=0$. Without loss of generality we may assume $\left({\varphi},1\right)_{\operatorname{L}^{2}(\Omega)}=0$, and according to Lemma 4.2 the function $\varphi$} satisfies
$\left\|{\varphi}\right\|_{\operatorname{PH}^{2}(\mathcal{G})}\leq C\left\|{\bf v}\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
where we used that ${\varphi}\in\operatorname{H}^{1}(\Omega)$ implies $\left(\left[{\varphi}\right],1\right)_{\operatorname{L}^{2}(\mathcal{I})}=0$. This proves $(i)$.
In order to prove $(ii)$, let ${\bf u}\in{\bf H}(\operatorname{curl},\Omega)$ be given and let ${\varphi}\in\operatorname{H}^{1}(\Omega)$ satisfy $\left({\varphi},1\right)_{\operatorname{L}^{2}{\Omega}}=0$ and
@@
for all $\xi\in\operatorname{H}^{1}(\Omega)$. Note that the Lax-Milgram lemma asserts the existence of $\varphi$}. By construction, we have $\left\|\nabla\varphi\right\|_{{\bf L}^{2}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and a Poincaré inequality shows $\left\|\varphi\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$. Furthermore, we observe that ${\bf v}:={\bf u}-\nabla\varphi\in{\bf H}(\operatorname{curl},\Omega)$ satisfies $\left(\boldsymbol{\nu}{\bf v},\nabla\xi\right)_{{\bf L}^{2}(\Omega)}=0$ for all $\xi\in{\bf H}^{1}(\Omega)$, and according to statement $(i)$ there holds $\left\|{\bf v}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf v}\right\|_{{\bf L}^{2}(\Omega)}=C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$. This proves $(ii)$.
If $\boldsymbol{\nu}$ is continuously differentiable on $\overline{\Omega}$, then the proof follows the lines, except for the fact that the equation $\operatorname{div}\boldsymbol{\nu}\nabla{\varphi}=\operatorname{div}\boldsymbol{\nu}{\bf v}$ in $\Omega$ with $\boldsymbol{\nu}\nabla{\varphi}\cdot{\bf n}=-\boldsymbol{\nu}{\bf v}\cdot{\bf n}$ on $\Gamma$ is a standard Neumann problem for $\varphi$}. ∎
@@
${a({\bf z},{\bf w}):=}\left(\boldsymbol{\nu}^{-1}\operatorname{curl}{\bf z},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=\left({\bf u},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}$
for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$. Note that according to Lemma 4.5, the sesquilinear for $a(\cdot,\cdot)$ is coercive on the space of vector fields ${\bf v}\in{\bf H}(\operatorname{curl},\Omega)$ that are orthogonal to all gradient fields, hence the Lax-Milgram lemma asserts the existence of $\bf z$}. Moreover, we have $\left\|\operatorname{curl}{\bf z}\right\|_{{\bf L}^{2}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ and exploiting Lemma 4.5 leads to